OpenStudy (anonymous):

3y/y(2y-1) + 2/2y-1 simplify lowest terms factored

OpenStudy (anonymous):

=3/(2y-1)+2/(2y-1) =5/(2y-1)

OpenStudy (anonymous):

so the final answer?

OpenStudy (anonymous):

that is the final answer

OpenStudy (anonymous):

5/2y-1??